See e-mail from Michael L.:

Dear Friends of ARGOS,

Yesterday (2/18/15) and today (2/19/15), ARGOS activities included the following:

Installation and test of the DX LM1 actuators by Matthias D., Matthias O. and Franz S.
installation of the DX limit switches delayed until tomorrow due to late delivery from DigiKey
originally scheduled for overnight delivery (2/18), severe weather in Tennessee caused major delay
order (finally) received at basecamp at 3:15 pm and will be delivered to summit in the morning
working with telescope crew, MPE and Arcetri to allow 1 hour at horizon for installation tomorrow
departure of Matthias O. and Franz S. today at 3 pm

Installation of ethernet, thermocouple and fiberoptic cables in the secondary swing arm cable trays by Julian Z.

Arrival of the Arcetri team on summit today; Lorenzo, Marco and Tomasso

Power up DX LGS WFS rack by Lorenzo, Marco, Tomasso and Michael
initialize system
perform flat field
measurement of pockets cell extinction
review of guis and procedures


Tomorrow morning we will start at horizon and complete installation of the LM1 limit switches. In addition, we will perform a full functional test of the LM1’s electrical, mechanical and software to verify operation and reliability. Next, we will revert to zenith and commence lgs wfs tests using the ARGOS swing arm. The goal is to acquire the light source, obtain good pointing/alignment of the spots and close the loop. LUCI2 will be used for acquisition (if required). Also, we will check tip/tilt using the pyramid wavefront sensor.

Gilles will also be joining our team tomorrow.

Tomorrow will be a full day!

All is well at the LBTO.

-- MarcoBonaglia - 20 Feb 2015
Topic revision: r1 - 20 Feb 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback